how to solve equations with variables on both sides of wqual sides? 0-13-4354422-2

Answer by checkley71(8403)

EXAMPLE: SAME VARIABLE ON BOTH SIDES
3X+4=12-X ADD X TO BOTH SIDES AND SUBTRACT 4 FROM BOTH SIDES
3X+X+4-4=12-4-X+X
4X+0=8-0X
4X=8
X=8/4
X=2
PROOF
3*2+4=12-2
6=4=10
10=10
HOPE THIS HELPS